OpenStudy (anonymous):

Find the derivative of the function: y= x^3 / x-1

OpenStudy (anonymous):

y = x * 7^(-3x) y ' = x * ( 7^(-3x) * ln(7) * -3 ) + 7^(-3x) * 1 y ' = -3x * 7^(-3x) * ln(7) + 7^(-3x) y ' = 7^(-3x) ( -3x * ln(7) + 1 )

OpenStudy (anonymous):

i think... i hope that helps.. :)

OpenStudy (anonymous):

Answer is y' =2x^3 - 3x^2/ (x-1)^2 but I just don't know how they got that.